1. Myungsung Co., Ltd. face value of 1,000,000 won, face rate 10% (interest payment date is the end of each year), 10,000 The 3-year bond was issued on January 1, 2015. The market interest rate applied at the time of issuance It is 7%. [When calculating the present value, you can calculate it directly or use the coefficient tables below. Some singular differences may be ignored depending on the mountain method.]


[Reference 1] Current value coefficient table

Years, interest rate

6%

7%

8%

9%

10%

1

0.943

0.935

0.926

0.917

0.909

2

0.890

0.873

0.857

0.842

0.826

3

0.840

0.816

0.794

0.772

0.751

4

0.792

0.763

0.735

0.708

0.683

5

0.747

0.713

0.681

0.650

0.621

[Reference 2] Table of present value of pension

Years, interest rate

6%

7%

8%

9%

10%

1

0.943

0.935

0.926

0.917

0.909

2

1.833

1.808

1.783

1.759

1.736

3

2.673

2.624

2.577

2.531

2.487

4

3.465

3.387

3.312

3.240

3.170

5

4.212

4.100

3.993

3.890

3.791

  • Question

  1. Please provide the accounting treatment to be performed in connection with the issuance of bonds on January 1, 2015.
  2. Calculate the amount of interest expense to be recognized in 2016 by Myungsung Co., Ltd. and provide the accounting treatment to be performed in connection with the calculation of interest expense on December 31, 2016.
  3. On January 1, 2017, Myeongsung Co., Ltd. repaid the total amount of bonds early and paid the creditors 1,100,000 won. Please provide the accounting to be performed in this regard
Answer #1
Amount PV Factor PV
Interest Payments            100,000 2.624              262,400
Face Value at maturity         1,000,000 0.816              816,000
Total           1,078,400
Date Account Titles Debit Credit
Jan-01 Cash         1,078,400
2015       Bonds Payable          1,000,000
      Premium on Bonds Payable               78,400

Interest Expense for 2015 = 1078400 x 7% = 75,488
Interest Expense for 2016 = (1078400-100000+75488) x 7% = 73,772

Date Account Titles Debit Credit
Dec-31 Interest Expense              73,772
2016 Premium on Bonds Payable              26,228
      Cash             100,000
Date Account Titles Debit Credit
Jan-01 Bonds Payable         1,000,000
2017 Premium on Bonds Payable              27,660
Loss on retirement of bonds              72,340
      Cash          1,100,000

If you have any query, kindly comment with your query and please mark thumbs up.
Answer #1
Amount PV Factor PV
Interest Payments            100,000 2.624              262,400
Face Value at maturity         1,000,000 0.816              816,000
Total           1,078,400
Date Account Titles Debit Credit
Jan-01 Cash         1,078,400
2015       Bonds Payable          1,000,000
      Premium on Bonds Payable               78,400

Interest Expense for 2015 = 1078400 x 7% = 75,488
Interest Expense for 2016 = (1078400-100000+75488) x 7% = 73,772

Date Account Titles Debit Credit
Dec-31 Interest Expense              73,772
2016 Premium on Bonds Payable              26,228
      Cash             100,000
Date Account Titles Debit Credit
Jan-01 Bonds Payable         1,000,000
2017 Premium on Bonds Payable              27,660
Loss on retirement of bonds              72,340
      Cash          1,100,000

If you have any query, kindly comment with your query and please mark thumbs up.